Problem 13-20 Using CAPM [LO4] A stock has a beta of 1.50 and an expected return of 14 percent. A risk-free asset currently earns 2 percent. a. What is the expected return on a portfolio that is equally invested in the two assets? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) Expected return % b. If a portfolio of the two assets has a beta of .84, what are the portfolio weights? (Do not round intermediate calculations. Round your answers to 4 decimal places, e.g., 32.1616.) Weight of stock Risk-free weight c. If a portfolio of the two assets has an expected return of 9 percent, what is its beta? (Do not round intermediate calculations. Round your answer to 3 decimal places, e.g., 32.161.) Beta d. If a portfolio of the two assets has a beta of 3.00, what are the portfolio weights? (Do not round intermediate calculations. Negative amounts should be indicated by a minus sign. Enter your answers as a whole number.) Weight of stock Risk-free weight
Answer #1

w denotes the weight of Stock C in the portfolio and 1-w denotes the weight of risk free

1.
=0.5*14%+0.5*2%=8%

2.
w*1.5+(1-w)*0=0.84
=>w=0.56
Hence,
Stock 56%
Risk free 44%

3.
w*14%+(1-w)*2%=9%
=>w=7/12
Portfolio beta=7/12*1.5=0.875

4.
w*1.5+(1-w)*0=3
=>w=2
Hence,
Stock 200%
Risk free -100%